OpenStudy (anonymous):

A worm crawled 5 cm in 6 sec. How long will it take to crawl 55 cm if it continues at the same rate? Please help you guys

OpenStudy (cwrw238):

speed = distance / time = 5/6 cm/sec

jimthompson5910 (jim_thompson5910):

5 cm/6 sec = 55 cm/x sec 5/6 = 55/x 5x = 6*55 5x = 330 x = 330/5 x = 66 So it will take 66 seconds